OpenStudy (anonymous):

3.05]Which of the following is greatest? thirteen twenty-fifths zero point four five forty-three one hundredths twenty-nine fiftieths

OpenStudy (aum):

Make the denominator of each fraction 100 so it will be easier to compare: \(\Large \frac{13}{25}, ~~ 0.45, ~~ \frac{43}{100}, ~~ \frac{29}{50}\) can be rewritten as: \(\Large \frac{13*4}{25*4}, ~~ \frac{45}{100}, ~~ \frac{43}{100}, ~~ \frac{29*2}{50*2}\) = \(\Large \frac{52}{100}, ~~ \frac{45}{100}, ~~ \frac{43}{100}, ~~ \frac{58}{100}\) Since the denominators are all the same, the fraction with the highest numerator is the greatest number.